OpenStudy (anonymous):

what is the derivative of (x-1)^2(x-3)? I keep expanding it all out to get x^3-5x^2+6x-3 and then finding the derivative of that but I am wrong. So what is it?

OpenStudy (cwrw238):

you can use the product rule d(x-1)^2 d(x-3) ------- * (x-3) + ----- * (x-1)^2 dx dx

OpenStudy (anonymous):

uhh... so we would find the derivative of (x-3) and multiply it too (x-1)^2 then take the derivative of (x-1)^2 and multiply it by (x-3?

OpenStudy (cwrw238):

= 2(x - 1)(x - 3) + (x - 1)^2

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

so it would be (x-1)^2(1)+(x-3)(2(x-1))?

OpenStudy (cwrw238):

d(uv) / dx = u (dv/dx ) + v (du/dx) where u and v are functions of x

OpenStudy (anonymous):

so... did I get it right?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

yessssss thanks, I find these rules annoying yet helpful lol